OpenStudy (anonymous):

This is an easy question but, can someone help me? I just really can't understand how to find an answer If 3x^2-x-14 (with quadratic formula please)

OpenStudy (whpalmer4):

Okay, you need to find the values of \(x\) that make \[3x^2-x-14=0\]You can use the quadratic formula for that, which states that if you have a quadratic of the form\[ax^2+bx+c = 0,\ a\ne0\]the solutions are given by \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Compare your equation with the first equation I gave, determine the values of \(a,b,c\), then plug them into the formula and grind out the answer(s).

OpenStudy (anonymous):

Hint: x+2 is a factor.

OpenStudy (whpalmer4):

\(x+2\) is a factor, but that isn't much help if you're supposed to do the problem in a way that doesn't involve factoring...

OpenStudy (anonymous):